A 470 kg sailboat is pushed North by the wind with a force of 402 n and South by the current of the ocean with a force of 360 �. What is the acceleration of the boat?


1
Expert's answer
2021-02-10T18:23:05-0500
"a=\frac{F}{m}\\a=\frac{402+360}{470}=1.62\frac{m}{s^2}"